OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. Please show all of your work. -x -3y=9

OpenStudy (amistre64):

-x -3y=9 --------- 9 -x -3y --- + ---- = 1 9 9/3 x y ---- + ---- = 1 -9 -3 when x=0; y=-3 when y=0; x=-9

OpenStudy (amistre64):

or to match point notation it can be written like this: when x=0 ; y=-3 when x=-9 ; y=0